OpenStudy (anonymous):

Evalute log 2(x-2) + log 2 (x+5)=3

OpenStudy (anonymous):

@Zarkon

Nnesha (nnesha):

familiar with the log properties ?

OpenStudy (anonymous):

yes

Nnesha (nnesha):

okay so apply one of them at left side \[\huge\rm \log_2(x-2) \color{reD}{+} \log_2 (x+5)=3\] there is plus sign whichnoe one would you use ? quotient or product property ?

OpenStudy (anonymous):

product

Nnesha (nnesha):

yes \[\huge\rm log_b x + \log_b y = \log_b(x \times y)\] try to apply this on ur original question pretty sure you can do it!

OpenStudy (anonymous):

x^2+3x-10

Nnesha (nnesha):

yes \[\huge\rm log_2 (x^2+3x-10)=3\] now we have to convert log to exponential form

OpenStudy (anonymous):

x^2+3x=13

Nnesha (nnesha):

how did you get 13 ?

Nnesha (nnesha):

ohh well log_2 is still there \[\huge\rm log_b x + \log_b y = \color{ReD}{\log_b}(x \times y)\] \[\huge\rm\color{reD}{ log_2} (x^2+3x-10)=3\] log_b is a common factor

OpenStudy (anonymous):

now what do i do?

OpenStudy (unklerhaukus):

now, exponentiate both sides of the equation, with the base 2

OpenStudy (anonymous):

What does that mean?

Nnesha (nnesha):

 convert log to exponential form